# If It’s Not Enough, Make It So: Reducing Authentic Data Demand in Face Recognition through Synthetic Faces

### v1

@@ Abstract
Recent advances in deep face recognition have spurred a growing demand for large, diverse, and manually annotated face datasets. Acquiring authentic, high-quality data for face recognition has proven to be a challenge, primarily due to privacy concerns. Large face datasets are primarily sourced from web-based images, lacking explicit user consent. In this paper, we examine whether and how synthetic face data can be used to train effective face recognition models with reduced reliance on authentic images, thereby mitigating data collection concerns. First, we explored the performance gap among recent state-of-the-art face recognition models, trained with synthetic data only and authentic (scarce) data only. Then, we deepened our analysis by training a state-of-the-art backbone with various combinations of synthetic and authentic data, gaining insights into optimizing the limited use of the latter for verification accuracy. Finally, we assessed the effectiveness of data augmentation approaches on synthetic and authentic data, with the same goal in mind. Our results highlighted the effectiveness of FR trained on combined datasets, particularly when combined with appropriate augmentation techniques.
@@ I Introduction
Figure 1: Average verification accuracy on the testing dataset (y-axis) vs. number of authentic identities from CASIA-WebFace (WF) in the training dataset (x-axis). In all settings, the verification accuracy improved by increasing the number of authentic identities in the training dataset. Also, combining synthetic (black line, DCFace; green line, GANControl) with a limited subset of authentic datasets improved the FR performance, in comparison to the case where only a limited subset of authentic identities (red line) is used to train FR. The number of synthetic identities in the combined dataset experiments is fixed (10K).
Face databases include millions of images often gathered from the Internet without user consent, thereby giving rise to concerns on their legal and ethical use for FR development. Notably, the approval of the General Data Protection Regulation (GDPR) [1] by the European Union (EU) in 2018 has significantly amplified criticism concerning the privacy-related issues associated with the use of publicly available face datasets obtained without adequate consent. The GDPR explicitly affirms individuals’ ”right to be forgotten” and imposes more stringent regulations concerning the acquisition, distribution, and utilization of biometric data, making adherence to such regulations a key, intricate endeavor when applied to face datasets. Consequently, several databases frequently employed for FR training, including but not limited to [11, 3, 19], have been retracted by their creators, to prevent potential legal issues. These circumstances have raised doubts upon the continuity of FR research, as the availability of such fundamental resources has become uncertain.
Synthetic data has recently emerged as a promising alternative to genuine datasets for FR training [10], serving as a viable solution to address the prevailing legal and ethical concerns [34, 6, 8, 2, 28]. This trajectory of research has been significantly propelled by the advancement of Deep Generative Models (DGMs), which facilitate the generation of synthetic samples thanks to their ability to capture the probability distribution underlying a face dataset. This generative process can be constrained to various factors, including but not limited to age, facial expression, head pose, and lighting [14, 37, 38, 17]. The majority of DGM techniques employed for the creation of synthetic faces build upon Generative Adversarial Networks (GANs) [18, 14, 36, 37, 6] or Diffusion Models (DMs) [21, 33, 16].
@@ II Related Work
Very recently, IDiff-Face [5] and DCFace [27] adopted diffusion models to generate synthetic data for FR training, achieving SOTA verification accuracy for synthetic-based FR. Specifically, IDiff-Face included fizziness in the identity condition to induce variations in the generated data. DCFace proposed a two-stage generative framework. In the first stage, an image of a novel identity using an unconditional diffusion model is generated and an image style from the style bank is selected. In the second stage, the generated images and style are mixed using a dual conditional diffusion model. GANDiffFace [30] proposed to combine GAN and diffusion models to generate synthetic data for FR training that exhibit certain attributes such controllable demographic distribution.
Although these approaches [27, 5, 2, 6, 34] achieved relatively high verification accuracy when they were evaluated on authentic datasets, their performance is lower than the one obtained by FR models trained on authentic data [13]. To overcome this challenge, SFace [6] applied knowledge distillation on the embedding level between the model trained on the authentic data and the one trained on synthetic data. DigiFace [2] and DCFace [27] combined their synthetic datasets with a subset of the authentic CASIA-WebFace dataset [41], aiming at improving the verification accuracy of synthetic-based FR. Both methods demonstrated that synthetic-based verification accuracy can be improved using a limited set of authentic data. Wang et at. [40] proposed to combine synthetic data from SynFace [34] with a small and limited subset of authentic data. The work proposed to utilize the same backbone to extract features from synthetic and authentic identity and then learn multi-class classification through two classification layers for synthetic and authentic classes. However, these works [2, 27, 40] only utilized a subset of authentic data, with a fixed number of authentic identities, and they did not investigate the performance gain when different ratios of synthetic and authentic identities are combined in the FR training. Recently, Melzi et al. [31] organized a competition at WACV 2024, aiming to explore the use of synthetic data in FR training and to attract the development of solutions for synthetic-based FR. The challenge considered two main tasks, training FR only with synthetic data and training FR with both synthetic and authentic data.
@@ III-A Datasets Preparation
As the authentic data, we adopted the well-known CASIA-WebFace [41] and MS1MV2 [13] datasets. CASIA-WebFace consists of 0.5M images of 10K identities. MS1MV2 is the refined version of the MS-Celeb-1M dataset [19]. MS1MV2 contains 5.8M images of 85K identities. To provide comparable results across FR trained on authentic-only data and on synthetic and authentic combinations, we utilized a subset of 10K identities of MS1MV2 in the conducted experiments. This subset will be noted as M2-S and contains the first 10K identities from MS1MV2.
The synthetic datasets come from GAN-Control [37] and DCFace [27]. Each of these datasets contains 0.5M images of 10K identities, i.e., 50 images per identity. The GAN-Control dataset is generated using an attribute conditional GAN-based model [37] trained on the FFHQ dataset [24]. Each synthetic identity in the GAN-Control dataset is generated by fixing the identity condition and randomizing the attribute conditions, i.e., age, pose, illumination and expression. From now on, the GAN-Control dataset will be noted as GC.
@@
Figure 2: FR Training Paradigm Overview. Subsets of authentic and synthetic data are combined to form the training dataset.
@@ III-D Model Training
We relied on FR models having the well-known ResNet-50 architecture [20] as the backbone and CosFace [39] as the loss function; the latter is defined as:
@@
During the training, Stochastic Gradient Descent (SGD) is employed as an optimizer with an initial learning rate of 0.1. The learning rate is divided by 10 at 22, 30 and 40 training epochs [6]. The models have been trained for 40 epochs with a total batch size of 512.
@@ IV Experimental Results
Our experiments analyzed, at first, the performance gaps between FR models trained on authentic-only and synthetic-only data created with SOTA generative methods (Section IV-A). Then, we explored how combining different ratios of authentic and synthetic data can influence FR performance (Section IV-B) and whether the combination of (scarce) authentic and synthetic data can actually overcome the limitations of both data domains (Section IV-C). Finally, we explored the effects of the random augmentation technique, when applied to all training images and only to the synthetic ones (Section IV-D). Our experiments were conducted independently on the GC and DC synthetic datasets, without any data interchange and ensuring a clear separation between them. For each synthetic dataset, we incorporated varying percentages ($10\%,20\%,40\%,60\%,80\%,$ and $100\%$) of the authentic datasets’ identities.
IV-A Gap in FR training on Synthetic or Authentic data (RQ1)
In our initial analysis, we aimed to understand the performance disparities between FR models trained exclusively on authentic data and those exclusively trained on synthetic data generated through SOTA generative methods. This investigation sought to shed light on the impact of data source variation on model performance, while considering an equal number of identities across datasets.
To address this objective, Table I presents the verification accuracy for models trained exclusively on synthetic datasets (first rows group) and models trained with an increasing amount of authentic data from the CASIA-WebFace dataset only (second rows group). Consistent with previous research [2, 27], the model trained on the full authentic dataset showed the highest performance, with an average accuracy of $94.63\%$ across the five considered benchmarks. Notably, the model trained on DC data only (second row, average accuracy of $89.56\%$) surpassed the performance of the model trained on GC data only (first row, average accuracy of $79.38\%$).
IV-B Mixing Ratio Impact on Size-fixed Training Data (RQ2)
@@
TABLE I: Verification accuracy on size-fixed training sets (authentic dataset: WF). First rows group: accuracy on full synthetic datasets; Second rows group: accuracy on different subsets of authentic WF; Third and fourth rows group: accuracy on different subsets of authentic WF and synthetic GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding limited subsets of authentic data. Highest accuracy scores for each rows group in bold.
In our second analysis, considering a fixed total number of identities in the training dataset, we were interested in assessing whether using different combinations of authentic and synthetic data for training the considered models influences the resulting verification performance. Specifically, we set the total number of identities to $10\mathrm{K}$ and evaluated the FR performance using different ratios of synthetic and authentic identities in the training datasets.
The results presented in Table I (third and fourth rows groups) demonstrate that incorporating a limited number of authentic identities (e.g., $1\mathrm{K}$) resulted in notable improvements compared to models trained exclusively on synthetic identities only (first rows group). More precisely, by combining only $1\mathrm{K}$ authentic identities with $9\mathrm{K}$ identities from the DC dataset, we observed a performance improvement w.r.t. the respective fully synthetic baselines of $0.49\%$ on LFW, $3.90\%$ on AgeDB-30, $8.45\%$ on CFP-FP, $0.69\%$ on CA-LFW and $5.69\%$ on CP-LFW, with an overall average improvement of $3.68\%$ across datasets. Mixing the same $1\mathrm{K}$ authentic identities with the $9\mathrm{K}$ identities generated by the other method, GC, led to more noticeable improvements. We observed an improvement over the synthetic baselines of $5.31\%$ on LFW, $14.88\%$ on AgeDB-30, $20.35\%$ on CFP-FP, $7.94\%$ on CA-LFW and $17.66\%$ on CP-LFW, with an overall $12.60\%$ average improvement.
@@
Given the performance differences between synthetic and authentic baselines, a consistent improvement was observed on all benchmarks as the number of involved authentic identities increase. Mixing GC data with the authentic data (third rows group) steadily improves the average performances from $89.38\%$ ($9\mathrm{K}$ synthetic and $1\mathrm{K}$ authentic identities) to $94.46\%$ ($1\mathrm{K}$ synthetic and $9\mathrm{K}$ authentic identities). This trend persisted while combining DC data with the authentic data (fourth rows group), with an increase in the average accuracy from $92.86\%$ to $94.68\%$.
Models trained on only subsets of authentic datasets, such as using $5\mathrm{K}$ authentic identities and $5\mathrm{K}$ synthetic ones, demonstrated highly competitive results (average accuracy of $99.30\%$ for GC and $94.28\%$ for DC) compared to models trained on the entire authentic dataset (average accuracy of $94.63\%$). Notably, certain combinations that involved DC identities even achieved superior performances than the full authentic-only baseline. Specifically, these observations were made for the settings with $2\mathrm{K}$ to $3\mathrm{K}$ synthetic identities for AgeDB-30, $1\mathrm{K}$ to $2\mathrm{K}$ synthetic identities for CFP-FP, and $1\mathrm{K}$ to $4\mathrm{K}$ synthetic identities for CP-LFW.
IV-C Performance Analysis as Authentic Data Increases (RQ3)
In our third analysis, we explored whether a training dataset with a large, fixed number of synthetic identities and a progressively growing number of authentic ones could achieve accuracy levels, at the very least, comparable to a larger dataset consisting solely of authentic identities, and we investigated the minimal ratio of authentic data required. To this end, Table II and III present the FR verification accuracy achieved by models trained on a large, fixed number of synthetic identities ($10\mathrm{K}$) and increasing amounts of authentic identities ($1\mathrm{K}$, $2\mathrm{K}$, $4\mathrm{K}$, $6\mathrm{K}$, $8\mathrm{K}$, $10\mathrm{K}$) from the WF and MS-2 datasets, respectively.
In both tables, we can observe that the accuracy becomes consistently higher by increasing the number of authentic identities in the training dataset, starting from a fixed number of synthetic identities from either GC or DC ($10\mathrm{K}$). This observation is evident when we compare the average accuracy estimates across rows groups in each table. By combining WF subsets with $10\mathrm{K}$ GC identities, models achieved an average accuracy ranging from $89.94\%$ ($1\mathrm{K}$ authentic identities) to $94.73\%$ ($10\mathrm{K}$ authentic identities). The same trend was observable also while combining WF subsets with DC identities, which led to an accuracy ranging from $93.09\%$ ($1\mathrm{K}$ authentic identities) to $95.06\%$ ($10\mathrm{K}$ authentic identities). However, it is worth noting that the accuracy obtained by combining the full $10\mathrm{K}$-identity synthetic dataset with $10\mathrm{K}$ authentic identities ($95.06\%$) were slightly worse than the ones obtained by combining it with $8\mathrm{K}$ authentic identities ($95.08\%$) from WF. Mixing M2-S with both GC and DC identities led to the same type of observation, reporting an accuracy ranging from $83.66\%$ to $95.08\%$ and from $90.72\%$ to $92.75\%$, respectively. As also visually emerging from Figure 3, all the settings involving combined datasets achieved a higher accuracy than the one of their respective synthetic-only baseline (first two row), with an improvement often of about $30\%$, making further comparisons under another granularity unnecessary.
To gain deeper insights into the influence of a specific type of generated data (GC or DC), we narrowed our focus to comparing the verification accuracy between combined and authentic-only datasets, while maintaining an identical number of authentic identities. This comparison was conducted within the same rows group in both Table II and III.
@@ IV-C Performance Analysis as Authentic Data Increases (RQ3)
TABLE II: Verification accuracy on increasingly larger number of authentic identities in training sets (authentic dataset: WF). First group: accuracy of models trained on synthetic GC and DC datasets, each with 10K identities. First row in each group: accuracy of models trained by using different subsets of authentic WF only. Second and third rows in each group: accuracy of models obtained by combining different subsets of authentic WF with 10K synthetic identities from GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding a limited subset of authentic data. Moreover, combining synthetic with authentic datasets (e.g., GC (10K) $\cup$ WF (10K)) led to higher accuracy, in comparison to the case where FR is only trained on authentic WF dataset/subsets. For each rows group, the highest accuracy score is highlighted in bold.
@@
TABLE III: Verification accuracy on increasingly larger number of authentic identities in training sets (authentic dataset: M2-S). First group: accuracy of models trained on synthetic GC and DC datasets, each with 10K identities. First row in each group: accuracy of models trained by using different subsets of authentic M2-S only. Second and third rows in each group: accuracy of models obtained by combining different subsets of authentic M2-S with 10K synthetic identities from GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding a limited subset of authentic data. Moreover, combining synthetic with authentic datasets (e.g., GC (10K) $\cup$ M2-S (10K)) led to higher accuracy, in comparison to the case where FR is only trained on authentic M2-S dataset/subsets. For each rows group, the highest accuracy score is highlighted in bold.
@@
Figure 3: Verification accuracy on testing data (y-axis) vs. number of authentic identities in training data (x-axis). Top row refers to authentic data sampled from WF, bottom row to authentic data sampled from M2-S. The verification accuracy improved by increasing the number of authentic identities in the training dataset. Also, combining synthetic with a limited subset of authentic datasets (black and green lines) improved the FR performance in comparison to the case where only a limited subset of authentic identities (red lines) is used to train FR. The number of synthetic identities in the combined dataset experiments is fixed (10K). Plots in the top (bottom) row visually resemble the results reported in Table II (Table III).
@@
Our results show, for small numbers of authentic identities, the improvement gained from adding the large synthetic dataset is more noticeable; however, it becomes less apparent when the number of authentic identities is higher. Specifically, training the models on DC (GC) data combined with authentic identities subsets from WF led, on average, to $10.83\%$ ($6.49\%$), $5.02\%$ ($3.30\%$), $2.81\%$ ($1.57\%$), $0.85\%$ ($-0.10\%$), $1.26\%$ ($0.69\%$) and $0.45\%$ ($0.11\%$) improvements w.r.t. the baseline trained only on authentic data. Each comparison was made by taking into consideration an authentic training subset involving 1K, 2K, 4K, 6K, 8K and 10K authentic identities, respectively. It also emerges that, under a given amount of authentic data, the impact of DC data is more evident than that of GC data.
@@
Following the same protocol, the training of FR models on the full DC (GC) dataset combined with authentic identity subsets from M2-S led to improvements in accuracy by $13.78\%$ ($4.93\%$), $9.41\%$ ($3.96\%$), $1.36\%$ ($1.69\%$), $-0.86\%$ ($0.51\%$), $-0.59\%$ ($1.33\%$), and $-1.96\%$ ($0.42\%$). These results, for small numbers of authentic identities, confirm that the inclusion of the full synthetic dataset results in a substantial increase of accuracy and that DC identities are more effective that those from GC. However, differently from what we observed for WF authentic data, the performance on DC identities combined with 6K or more authentic identities from M2-S not only achieved lower performances compared to combining with GC identities, but also adversely affected performance with respect to using the authentic data only. This observation sheds a light on the importance of devising future work about methods for creating synthetic data functional to improving model accuracy, in an end-to-end manner.
IV-D Augmentation Impact on Combined Training Data (RQ4)
In our final analysis, we investigated the impact of data augmentation, as specified in Section III-C, on verification accuracy concerning combined datasets. Our experiments involved applying data augmentation to two distinct subsets of the training dataset: exclusively on synthetic identities and on the entire combined dataset. We compared the results from both scenarios with the ones that emerged from the previous research question, which will constitute the baselines in this subsection. To address this, Table IV and V illustrate the FR verification accuracy achieved by models trained on a fixed number of synthetic identities ($10\mathrm{K}$, from GC and DC respectively) combined with increasing amounts of authentic identities ($1\mathrm{K}$, $2\mathrm{K}$, $4\mathrm{K}$, $6\mathrm{K}$, $8\mathrm{K}$, $10\mathrm{K}$) from WF. The augmentation was applied to both authentic and synthetic subsets or only to the latter.

### v4

@@ Abstract
Recent advances in deep face recognition have spurred a growing demand for large, diverse, and manually annotated face datasets. Acquiring authentic, high-quality data for face recognition has proven to be a challenge, primarily due to privacy concerns. Large face datasets are primarily sourced from web-based images, lacking explicit user consent. In this paper, we examine whether and how synthetic face data can be used to train effective face recognition models with reduced reliance on authentic images, and thus mitigating large authentic data collection concerns. First, we explored the performance gap among recent state-of-the-art face recognition models, trained only on synthetic data or authentic data. Then, we deepened our analysis by training a state-of-the-art backbone with various combinations of synthetic and authentic data, gaining insights into optimizing the limited use of the latter for verification accuracy. Finally, we assessed the effectiveness of data augmentation approaches on synthetic and authentic data, with the same goal in mind. Our results highlighted the effectiveness of FR trained on combined datasets, particularly when combined with appropriate augmentation techniques.
@@ I Introduction
Figure 1: Average verification accuracy on the testing dataset (y-axis) vs. number of authentic identities from CASIA-WebFace (WF) [44] in the training dataset (x-axis). In all settings, the verification accuracy improved by increasing the number of authentic identities in the training dataset. Also, combining synthetic (black line: DCFace [28], green line: ExFaceGAN(GANControl) [7]) with a limited subset of authentic data improved the FR performance, in comparison to the case where only a limited subset of authentic identities (red line) is used to train FR. The number of synthetic identities in the combined dataset experiments is fixed (10K).
Face databases include millions of images often gathered from the Internet without user consent, thereby giving rise to concerns about their legal and ethical use for FR development. Notably, the approval of the General Data Protection Regulation (GDPR) [1] by the European Union (EU) in 2018 has significantly amplified criticism concerning the privacy-related issues associated with the use of publicly available face datasets obtained without adequate consent. The GDPR explicitly affirms individuals’ ”right to be forgotten” and imposes more stringent regulations concerning the acquisition, distribution, and utilization of biometric data, making adherence to such regulations a key, intricate endeavor when applied to face datasets. Consequently, several databases frequently employed for FR training, including but not limited to [11, 3, 20], have been retracted by their creators, to prevent potential legal issues. These circumstances have raised doubts about the continuity of FR research, as the availability of such fundamental resources has become uncertain.
Synthetic data has recently emerged as a promising alternative to genuine datasets for FR training [10], serving as a viable solution to address the prevailing legal and ethical concerns [36, 6, 8, 2, 29]. This research trajectory has been significantly propelled by the advancement of Deep Generative Models (DGMs), which facilitate the generation of synthetic samples thanks to their ability to capture the probability distribution underlying a face dataset. This generative process can be constrained to various factors, including but not limited to age, facial expression, head pose, and lighting [15, 40, 41, 18]. The majority of DGM techniques employed for the creation of synthetic faces build upon Generative Adversarial Networks (GANs) [19, 15, 39, 40, 6] or Diffusion Models (DMs) [22, 35, 17].
@@ II Related Work
Very recently, IDiff-Face [5] and DCFace [28] adopted diffusion models to generate synthetic data for FR training, achieving SOTA verification accuracy for synthetic-based FR. Specifically, IDiff-Face included fizziness in the identity condition to induce variations in the generated data. DCFace proposed a two-stage generative framework. In the first stage, an image of a novel identity using an unconditional diffusion model is generated and an image style from the style bank is selected. In the second stage, the generated images and style are mixed using a dual conditional diffusion model. GANDiffFace [31] proposed to combine GAN and diffusion models to generate synthetic data for FR training that exhibit certain attributes such as controllable demographic distribution.
Although these approaches [28, 5, 2, 6, 36] achieved relatively high verification accuracy when they were evaluated on authentic datasets, their performance is lower than the one obtained by FR models trained on authentic data [14]. To overcome this challenge, SFace [6] applied knowledge distillation on the embedding level between the model trained on the authentic data and the one trained on synthetic data. DigiFace [2] and DCFace [28] combined their synthetic datasets with a subset of the authentic CASIA-WebFace dataset [44], aiming at improving the verification accuracy of synthetic-based FR. Both methods demonstrated that synthetic-based verification accuracy can be improved using a limited set of authentic data. Wang et at. [43] proposed to combine synthetic data from SynFace [36] with a small and limited subset of authentic data. The work proposed to utilize the same backbone to extract features from synthetic and authentic identity and then learn multi-class classification through two classification layers for synthetic and authentic classes. However, these works [2, 28, 43] only utilized a subset of authentic data, with a fixed number of authentic identities, and they did not investigate the performance gain when different ratios of synthetic and authentic identities combined in the FR training. Recently, several challenges and competitions have been organized in conjunction with top venues, aiming at promoting privacy-friendly synthetic-based FR development. FRCSyn competition [32] was organized at WACV 2024, aiming to explore the use of synthetic data in FR training and to attract the development of solutions for synthetic-based FR. The challenge considered two main tasks, training FR only with synthetic data and training FR with both synthetic and authentic data. The achieved results of the top-performing solutions from FRCSyn [32] competition are further investigated and reported in [33]. A second edition of FRCSyn [13] was organized in conjunction with CVPR 2024. Also, very recently SDFR [38] competition was organized in conjunction with FG 2024, attracting developing and promoting solutions for synthetic-based FR.
@@ III-A Datasets Preparation
As the authentic data, we adopted the well-known CASIA-WebFace [44] and MS1MV2 [14] datasets. CASIA-WebFace [44] consists of 0.5M images of 10K identities. MS1MV2 is the refined version of the MS-Celeb-1M dataset [20]. MS1MV2 contains 5.8M images of 85K identities. To provide comparable results across FR trained on authentic-only data and on synthetic and authentic combinations, we utilized a subset of 10K identities of MS1MV2 in the conducted experiments. This subset will be noted as M2-S and contains the first 10K identities from MS1MV2.
The synthetic datasets come from ExFaceGAN [7] and DCFace [28]. Each of these datasets contains 0.5M images of 10K identities, i.e., 50 images per identity. The ExFaceGAN dataset is generated using an identity disentanglement approach [7] on pretrained GAN-Control [40] which is trained on the FFHQ dataset [25]. From now on, the ExFaceGAN(GAN-Control) dataset will be noted as GC.
@@
Figure 2: FR Training Paradigm Overview. Subsets of authentic and synthetic data are combined to form the training dataset. During the training phase, only synthetic data is augmented with RandAugment, as well discussed in Section IV. the utilized network architecture in all settings is ResNet50 [21] trained with CosFace [42] loss.
@@ III-D Model Training
We relied on FR models having the well-known ResNet50 architecture [21] as the backbone and CosFace [42] as the loss function; the latter is defined as:
@@
During the training, Stochastic Gradient Descent (SGD) is employed as an optimizer with an initial learning rate of 0.1. The learning rate is divided by 10 at 22, 30, and 40 training epochs [6]. The models have been trained for 40 epochs with a total batch size of 512.
@@ IV Experimental Results
Our experiments analyzed, at first, the performance gaps between FR models trained on authentic-only and synthetic-only data created with SOTA generative methods (Section IV-A). Then, we explored how combining different ratios of authentic and synthetic data can influence FR performance (Section IV-B) and whether the combination of authentic and synthetic data can actually overcome the limitations of both data domains (Section IV-C). Finally, we investigate the impact of augmentation on synthetic FR performance (Section IV-D). Our experiments were conducted on the GC and DC synthetic datasets.
IV-A Performance gap between FR trained on an authentic dataset and ones trained on synthetic dataset
Table I presents the verification accuracies on five benchmarks achieved by models trained on GC, DC and WF datasets. The table also presents the achieved accuracies of FR models trained on different subsets of authentic WF. As expected, the model trained on authentic WF (10K) outperformed the models trained on synthetic GC (10K) and DC (10K) [28]. The average verification accuracies achieved by the model trained on WF (10K) is 94.63% and the achieved ones by the models trained on GC and DC are 79.38% and 89.56%, respectively. It can be also observed that the model trained on a subset of 3K identities of WF (average accuracies of 90.83%) outperformed models trained on 10K identities of synthetic DC and GC datasets. This large performance gap can be reduced by combining synthetic with a small set of authentic identities as will be presented in the next section.
IV-B Impact of combining synthetic data with a small subset of authentic data on FR performance
@@
TABLE I: Verification accuracy on size-fixed training sets (authentic dataset: WF). First rows group: accuracy on full synthetic datasets. Second rows group: accuracy on different subsets of authentic WF. Third and fourth rows group: accuracy on different subsets of authentic WF and synthetic GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding limited subsets of authentic data. Highest accuracy scores for each row group in bold.
The third and fourth groups of results in Table I present the achieved FR verification accuracies of FR models trained on different datasets that are built on by combining subsets of authentic and synthetic identities. The total number of identities in each case is fixed to be equal to 10K, while the ratio of synthetic and authentic identities varies.
Combining only $1\mathrm{K}$ authentic identities with $9\mathrm{K}$ identities from DC enhanced the FR model accuracy w.r.t. the DC synthetic baseline of $0.49\%$ on LFW, $3.90\%$ on AgeDB-30, $8.45\%$ on CFP-FP, $0.69\%$ on CA-LFW and $5.69\%$ on CP-LFW, with an overall average improvement of $3.68\%$ across datasets. Combining the same $1\mathrm{K}$ authentic identities with $9\mathrm{K}$ synthetic identities from GC resulted in even more noticeable improvements over the respective baseline: $5.31\%$ on LFW, $14.88\%$ on AgeDB-30, $20.35\%$ on CFP-FP, $7.94\%$ on CA-LFW and $17.66\%$ on CP-LFW, with an overall $12.60\%$ average improvement.
@@
As expected, a consistent improvement was observed on all benchmarks as the ratio of involved authentic identities increased. Combining subsets of identities from GC and WF (third group of rows) steadily improved the average verification accuracy from $89.38\%$ ($9\mathrm{K}$ synthetic and $1\mathrm{K}$ authentic identities) to $94.46\%$ ($1\mathrm{K}$ synthetic and $9\mathrm{K}$ authentic identities). The same trend was also noted while combining DC subsets with the same authentic identities (fourth group of rows), with an increase in the average accuracy from $92.86\%$ to $94.68\%$.
IV-C Impact of combining fixed numbers of synthetic identities with different subsets of authentic identities on FR performance
Tables II and III present the FR verification accuracy achieved by models trained on a fixed number of synthetic identities ($10\mathrm{K}$) combined with a subset of authentic identities ($1\mathrm{K}$, $2\mathrm{K}$, $4\mathrm{K}$, $6\mathrm{K}$, $8\mathrm{K}$, or $10\mathrm{K}$) from the WF and MS-2 authentic datasets, respectively.
Training FR models on subsets of authentic identities combined with synthetic data consistently improved the verification accuracy as the number of authentic identities in the training dataset increased. Models trained on a subset of 10K identities from WF combined with $10\mathrm{K}$ identities from GC achieved average verification accuracy improvements from $89.94\%$ ($1\mathrm{K}$ authentic identities) to $94.73\%$ ($10\mathrm{K}$ authentic identities). Similarly, training on a subset of 10K identities from WF combined with DC improved the average verification accuracy from $93.09\%$ ($1\mathrm{K}$ authentic identities) to $95.06\%$ ($10\mathrm{K}$ authentic identities). Notably, the performance obtained by training the model on DC combined with $10\mathrm{K}$ authentic identities ($95.06\%$) was slightly worse than the one obtained by combining it with $8\mathrm{K}$ authentic identities ($95.08\%$) from WF.
Training FR models on M2-S combined with either GC or DC led to similar observations, reporting verification accuracy improvements from $83.66\%$ to $95.08\%$ and from $90.72\%$ to $92.75\%$, respectively. As visually emerging from Figure 3, all the settings in which the models were trained on combined datasets achieved higher verification accuracy than their respective synthetic baseline, with best improvements of about $30\%$.
@@
TABLE II: Verification accuracy of models trained on synthetic datasets combined with different subsets authentic WF. First group: accuracy of models trained on synthetic GC and DC datasets, each with 10K identities. The first row in each group: accuracy of models trained by using different subsets of authentic WF only. Second and third rows in each group: accuracy of models obtained by combining different subsets of authentic WF with 10K synthetic identities from GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding a limited subset of authentic data. Moreover, combining synthetic with authentic datasets (e.g., GC (10K) $\cup$ WF (10K)) led to higher accuracy, in comparison to the case where FR is only trained on authentic WF dataset/subsets. For each row group, the highest accuracy score is highlighted in bold.
@@
TABLE III: Verification accuracy of models trained on synthetic datasets combined with different subsets authentic M2-S. The first group: accuracy of models trained on synthetic GC and DC datasets, each with 10K identities. The first row in each group: accuracy of models trained by using different subsets of authentic M2-S only. Second and third rows in each group: accuracy of models obtained by combining different subsets of authentic M2-S with 10K synthetic identities from GC and DC datasets, respectively. It can be noticed that the accuracy of synthetic-based models improved by adding a limited subset of authentic data. Moreover, combining synthetic with authentic datasets (e.g., GC (10K) $\cup$ M2-S (10K)) led to higher accuracy, in comparison to the case where FR is only trained on authentic M2-S dataset/subsets. For each row group, the highest accuracy score is highlighted in bold.
@@
Figure 3: Average verification accuracy on testing data (y-axis) vs. number of authentic identities in training data (x-axis). The top row figures refer to authentic data sampled from WF, second row figures refer to authentic data sampled from M2-S. The verification accuracy improved by increasing the number of authentic identities in the training dataset. Also, combining synthetic with a limited subset of authentic datasets (black and green lines) improved the FR performance in comparison to the case where only a limited subset of authentic identities (red lines) is used to train FR. The number of synthetic identities in the combined dataset experiments is fixed (10K). The results in these plots correspond to the ones reported in Table II Table III.
As visually illustrated in Figure 3, increasing the number of authentic identities did improve, as expected, the FR performances with the best performance achieved by using 10K authentic identities, i.e. utilizing the whole authentic dataset. However, similar performances, or even higher, can be achieved by utilizing a small set of authentic identities, e.g. 4K authentic identities, combined with 10K synthetic identities from either GC or DC.
@@
Training the models on GC combined with authentic identity subsets from WF led to $6.49\%$, $3.30\%$, $1.57\%$, $-0.10\%$, $0.69\%$ and $0.11\%$ improvements in terms of average verification accuracy. Similarly, training on DC combined WF subsets led to $10.83\%$, $5.02\%$, $2.81\%$, $0.85\%$, $1.26\%$ and $0.45\%$ improvements in terms of average verification accuracy. Each comparison was made by taking into consideration an authentic subset involving 1K, 2K, 4K, 6K, 8K, and 10K identities, respectively. It also emerges that, under a given amount of authentic identities, the impact of DC is more evident than that of GC.
@@
Following the same protocol, training FR models on DC (GC) combined with authentic identity subsets from M2-S led to average verification accuracy improvements of $13.78\%$ ($4.93\%$), $9.41\%$ ($3.96\%$), $1.36\%$ ($1.69\%$), $-0.86\%$ ($0.51\%$), $-0.59\%$ ($1.33\%$), and $-1.96\%$ ($0.42\%$). These results further confirm that training on a large synthetic dataset combined with small authentic identity subsets results in a substantial increase in verification accuracy and that DC identities are more effective than those from GC. However, differently from what we observed for WF, the performance achieved by training a FR model on DC combined with 6K or more authentic identities from M2-S not only achieved lower performances compared to combining with GC identities but also adversely affected performance with respect to training only on authentic data. This observation sheds light on the importance of devising future work about methods for creating synthetic data functional to improve model accuracy, in an end-to-end manner.
IV-D Impact of data augmentation
Tables IV and V present the achieved verification accuracies by FR models trained on different subsets of authentic and synthetic data and by augmenting the training dataset with RandAug presented in Section III-C.